OpenStudy (anonymous):

Find the limit as x->0 (sin 2x)/(4x)

OpenStudy (campbell_st):

well to use \[\lim_{x \rightarrow 0} \frac{\sin(x)}{x} = 1\] you need the x in the numberator and denominator to be he same so rewrite it as \[\lim_{x \rightarrow 0} \frac{1}{2} \times \frac{\sin(2x)}{2x} = \frac{1}{2} \lim_{x \rightarrow 0} \frac{\sin(2x)}{2x}\]

OpenStudy (anonymous):

how did 4x turn into 2x?? .__.

OpenStudy (campbell_st):

well 1/2 * 1/2x = 1/(2*2x) = 1/4x

OpenStudy (anonymous):

why did you use 1/2?

OpenStudy (anonymous):

lol i'm lost :(

OpenStudy (anonymous):

why did you use 1/2?